OpenStudy (anonymous):

The distance between A (-5, 3) and B(5, 3) is. *3 *6 *10

OpenStudy (ivettef365):

use distance formula

hartnn (hartnn):

Distance between points (x1,y1) and (x2,y2) is \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\) does this help ??

OpenStudy (anonymous):

Its not workingin my calculator..

hartnn (hartnn):

ok, do it manually 5-(-5) = ...?

OpenStudy (dls):

hadd he hogayi :|

OpenStudy (anonymous):

10?

hartnn (hartnn):

yes, correct. and 3-3 =.... ?

OpenStudy (anonymous):

0

hartnn (hartnn):

now, \(\sqrt{10^2+0^2}=\sqrt{10^2}=....?\)

OpenStudy (dls):

genius :O

OpenStudy (anonymous):

10?

hartnn (hartnn):

correct! thats your distance...

OpenStudy (anonymous):

Thanks!

OpenStudy (dls):

don't forget the "units"

hartnn (hartnn):

welcome ^_^